OpenStudy (anonymous):

Which of the following statements best describes the effect of replacing the graph of y = f(x) with the graph of y = f(x) + 9? (3 points)

OpenStudy (anonymous):

The graph of y = f(x) will shift up 9 units. The graph of y = f(x) will shift down 9 units. The graph of y = f(x) will shift left 9 units. The graph of y = f(x) will shift right 9 units.

OpenStudy (anonymous):

Someone please help me, i am so lost

OpenStudy (jingle535):

Hey, when you add 9 that's adding 9 to the y intercept so you do you have idea what that would do? (I know the answer I want to try to make sure you understand it :))

OpenStudy (anonymous):

Shift up 9 units?

OpenStudy (jingle535):

Yes, Amazing job!

OpenStudy (anonymous):

Yaye thank you!

OpenStudy (anonymous):

Could you help me with one more please?

OpenStudy (anonymous):

The functions f(x) and g(x) are described below: f(x) = 3x - 7 g(x) = 3x + 8 The graph of g(x) is obtained by shifting up the graph of f(x) by _____ units. (3 points)

OpenStudy (jingle535):

Yes I can, now this is another one with the y intercept so how does the -7 change to be 8? And fyi I gave you a medal because I helped you answer your own question :)

OpenStudy (anonymous):

Would you subract -7 from 8?

OpenStudy (jingle535):

yes :)

OpenStudy (anonymous):

so would the answer be -15?

OpenStudy (anonymous):

wait im a little confused

OpenStudy (jingle535):

Sooooo close you just messed up on the addition, when you subtract a negative number it adds the negative number. Here is an example 3 - -4= 7

OpenStudy (anonymous):

i did -7 - 8 and i got -15

OpenStudy (jingle535):

Think of it this way, how many numbers do you need to get from -7 to 8?

OpenStudy (anonymous):

15

OpenStudy (jingle535):

That's your answer :)

OpenStudy (anonymous):

Oh haha thanks :)

OpenStudy (anonymous):

oh thats right because you cant move negative units away from something XD

OpenStudy (jingle535):

No problem :)

OpenStudy (anonymous):

An airplane ascends at the rate of 40 feet per minute. The function below shows the height of the airplane from the ground f(t) after t minutes: f(t) = 40t + 520 Which statement best describes the y-intercept of the graph of the function? (3 points)

OpenStudy (anonymous):

one more??? pweez

OpenStudy (anonymous):

It is equal to 520 and represents the height of the airplane above the ground, in feet, after 1 minute. It is equal to 560 and represents the height of the airplane above the ground, in feet, after 1 minute. It is equal to 40 and represents the height of the airplane above the ground, in feet, before it starts ascending. It is equal to 520 and represents the height of the airplane above the ground, in feet, before it starts ascending.